### v3

# One latent to fit them all: a unified representation of baryonic feedback on matter distribution

Abstract
Accurate and parsimonious quantification of baryonic feedback on matter distribution is of crucial importance for understanding both cosmology and galaxy formation from observational data. This is, however, challenging given the large discrepancy among different models of galaxy formation simulations, and their distinct subgrid physics parameterizations. Using 5,072 simulations from 4 different models covering broad ranges in their parameter spaces, we find a unified 2D latent representation. Compared to the simulations and other phenomenological models, our representation is independent of both time and cosmology, much lower-dimensional, and disentangled in its impacts on the matter power spectra. The common latent space facilitates the comparison of parameter spaces of different models and is readily interpretable by correlation with each. The two latent dimensions provide a complementary representation of baryonic effects, linking black hole and supernova feedback to distinct and interpretable impacts on both the matter power spectrum, and field, level. Our approach enables developing robust and economical analytic models for optimal gain of physical information from data, and is generalizable to other fields with significant modeling uncertainty.
[…]
journal: ApJLsoftware: PyTorch (A. Paszke et al., 2019), Optuna (T. Akiba et al., 2019), NumPy (C. R. Harris et al., 2020), & matplotlib (J. D. Hunter, 2007).
show]shuruil3@illinois.edu
show]eelregit@gmail.com
I Introduction
Observational cosmology is entering a new era, marked by the rapid advancement of Stage-IV experiments, including Euclid (E. Collaboration et al., 2025), The Dark Energy Spectroscopic Instrument (DESI) (D. Collaboration et al., 2025), Legacy Survey of Space and Time (LSST) performed by Vera C. Rubin Observatory (Ž. Ivezić et al., 2019), and Nancy Grace Roman Space Telescope (Roman) (J. E. Schlieder et al., 2024). While most cosmological observables focus on the luminous baryonic matter in our Universe, weak gravitational lensing additionally has the unique strength in directly probing the total matter distribution (M. Kilbinger, 2015), that of both baryonic and dark matter. Optimal analyses of weak-lensing data require accurate modeling of baryonic physics on small scales, where galaxies form and feed back to the environment (T. Lu & Z. Haiman, 2021). However, exploiting small-scale weak-lensing data requires accurate baryonic physics models up to $k_{\mathrm{max}}\sim 0.3\ h/\mathrm{Mpc}$ (The LSST Dark Energy Science Collaboration et al., 2018), and quantitative agreement remains a significant obstacle, limiting our ability to make precise inferences and fully exploit the potential of forthcoming surveys (M. Vogelsberger et al., 2019; D. Copeland et al., 2018; H.-J. Huang et al., 2019). Therefore, an accurate model of baryonic effects at small scales is crucial to achieving precise cosmological measurements.
To study the non-linear structure formation and multi-scale galaxy formation, numerical simulations, especially hydrodynamical simulations, have become the essential tool. Hydrodynamical simulations model gas dynamics, star formation, and feedback mechanisms (T. Di Matteo et al., 2005; J. Salcido et al., 2023), with notable examples including IllustrisTNG, Astrid, Swift-EAGLE, and SIMBA (D. Nelson et al., 2021; R. Davé et al., 2019; S. Bird et al., 2022; R. A. Crain et al., 2015), offering a broad spectrum of feedback models and parameterizations. Building upon these developments, the Cosmology and Astrophysics with Machine Learning Simulations (CAMELS) project has recently generated an extensive set of cosmological simulations that systematically vary both cosmological and baryonic physics parameters across different feedback models (F. Villaescusa-Navarro et al., 2021; Y. Ni et al., 2023). This rich dataset enables a comprehensive investigation of baryonic feedback effects using advanced statistical and machine learning techniques.
Hydrodynamical simulations are computationally expensive, typically costing more than one order of magnitude than the $N$-body simulations that evolve total matter under only gravity. This high cost severely limits the ability to fully sample cosmological and astrophysical parameter spaces, creating a bottleneck for accurate modeling of baryonic feedback in large-scale structure analyses. Fortunately, baryonic feedback primarily alters the amplitude rather than the phase of Fourier modes in the matter density field, leaving the mode phases largely unchanged, even on small scales ($k\lesssim 10\ h\mathrm{Mpc}^{-1}$) across different hydrodynamical simulations (D. Sharma et al., 2024). Instead of evolving the hydrodynamical density field fully, one can focus on modeling how feedback modifies the amplitude of density fluctuations relative to a dark-matter-only baseline. A convenient way to encode this effect is through the squared transfer function ($T^{2}$) defined by the ratio of the hydrodynamic power spectra ($P_{\mathrm{hyd}}$) and dark-matter-only power spectra ($P_{\mathrm{dmo}}$):
$T^{2}(k,a)=\frac{P_{\mathrm{hyd}}(k,a)}{P_{\mathrm{dmo}}(k,a)},$ (1)
which enables efficient modeling of baryonic feedback at both the field level and the power-spectrum level of total matter distribution simply from the $N$-body runs, avoiding the cost of large hydrodynamic suites while retaining feedback effects.
With the growing availability of simulation datasets and advances in machine learning, several recent studies have modeled the transfer function $T^{2}$ via Emulation-based methods (D. Sharma et al., 2024; A. J. Zhou et al., 2025; M. Schaller et al., 2025), analytic formulations, (M. Schaller & J. Schaye, 2025; L. Kammerer et al., 2025) and physically motivated parameterizations (I. Medlock et al., 2025; M. P. van Daalen et al., 2019). Alternatively to hydrodynamical simulation, baryonification or baryon correction models apply parametrized prescriptions to displace matter distribution in gravity-only simulations to account for the influence of baryons (G. Aricò et al., 2020, 2021a, 2021b). Despite their successes, most approaches are limited to a single simulation model or a narrow parameter space, which can hinder generalization. Moreover, many introduce multiple, often redshift-dependent, parameters, leading to degeneracies that reduce information content and obscure physical interpretability. Collectively, these examples underscore the difficulty of integrating data from different simulation suites with a wide range of parameter choices and the challenge of developing a disentangled representation for baryonic feedback modeling. We also provide a summarized comparison in Appendix A.
With the development of machine learning, variational autoencoders (VAEs) (D. P. Kingma & M. Welling, 2022a; O. Rybkin et al., 2021) offer a way to learn latent representations adaptable to multiple datasets (L. Lucie-Smith et al., 2022; D. Piras & L. Lombriser, 2024). Furthermore, to improve the interpretability of the learned representations, C. P. Burgess et al. (2018) introduced the $\beta$-VAE, which encourages disentanglement in the latent space by increasing the weight on the KL divergence term. However, this often comes at the cost of reduced reconstruction accuracy and limited information capacity in the latent representation. R. T. Q. Chen et al. (2019) proposed the $\beta$-TCVAE, which incorporates additional hyperparameters to upweight the total correlation term in the loss function (see Section C.1) to explicitly disentangle the latent dimensions while preserving both reconstruction quality and information in the latent space.
In this paper, we propose a $\beta$-TCVAE–based representation-learning approach to reconstruct the transfer function $T^{2}$ and learn a latent representation of baryonic feedback across four CAMELS simulation suites over $0.356<k<10,h/\mathrm{Mpc}$ and $0\leq z\leq 2$. Based on these, we achieve the following major contributions:
  • Nonlinearity: The neural-network and variational capacity of our VAE model captures the non-linear dependence of baryonic feedbacks on galaxy formation physics, across thousands of simulations and multiple subgrid physics models without amplifying noise.
  • Independence: The latent space is trained to be independent of both redshift and cosmology, preventing duplicating the cosmological dependence and avoiding multiplying the latent dimensionality with number of redshift bins.
  • Parsimony: The latent space is only 2D, minimizing information loss when marginalized over in cosmological inference.
  • Interpretability: The two latent parameters exhibit distinct scale and time dependences, and associate with different feedback mechanisms by correlation with galaxy formation parameters.
This paper is organized as follows: In Section II.1, we introduce the datasets we used in this study. In Section II.2, we introduce the design of the VAE model and loss. We show the training strategy in Section II.3. In section III, we show the results of the reconstruction of spectrum ratio and the latent representation of baryonic feedback effects. In section IV, we discuss the results and the potential application of the latent representation. The appendices include more details, including, notably, a comparison with the principal component analysis (PCA) approach in Appendix I.
II Methods
II.1 Data
The Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) project comprises over 15,000 cosmological simulations, including 8,925 hydrodynamic and 6,136 N-body runs. In this paper, we utilize the first-generation simulations of CAMELS, each performed in a periodic box of $25~h^{-1}\,{\rm Mpc}$ per side, tracking the evolution of $256^{3}$ dark matter particles and $256^{3}$ initial fluid elements. This setup defines the baseline resolution for this simulation suite. Our analysis focuses on four specific hydrodynamic suites [^1] https://camels.readthedocs.io/en/latest/parameters.html , IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6, each characterized by variations in cosmological and astrophysical parameters, as well as distinct random seeds for initial conditions. We summarize the basic information of the four suites in Table 2 in Appendix B.
To assess the robustness of our results, we additionally incorporate the CAMELS Cosmic Variance (CV) set for benchmarking. This set comprises 27 simulations per suite, each sharing the same fiducial cosmological and astrophysical parameters but differing in the initial condition random seed. These simulations are commonly used to isolate the effects of cosmic variance. We use the CV counterparts of our main four suites: IllustrisTNG CV, Astrid CV, SIMBA CV, and Swift-EAGLE CV. For a detailed description of the above simulations, we refer the reader to F. Villaescusa-Navarro et al. (2021); Y. Ni et al. (2023).
For all suites, each hydrodynamic simulation is paired with a gravity-only simulation (which is also called a dark-matter-only simulation) with the same initial conditions. We compute the transfer function $T^{2}$ from the ratio of hydrodynamic to dark-matter-only total power spectra. Data are split into training, validation, and test sets in an 1/2: 7/16: 1/161/2\ {:}\ 7/16\ {:}\ 1/16 ratio to ensure robust values for the KL terms in validation loss (see Section D.1). From each simulation, we use three snapshots ($a\approx 1/3,1/2,1$), rebin the power spectra into 18 $k$-bins over $k\in[0.356,28.90]h\mathrm{Mpc}^{-1}$ and train the model with the $3\times 18$ dimensions spectrum ratio. Based on the requirement of LSST (The LSST Dark Energy Science Collaboration et al., 2018) and the noise level in the large-$k$ end, we limit our inference to $0.356<k<10\ h/\mathrm{Mpc}$.
II.2 Variational autoencoders
In this section, we breifly review Variational Autoencoders (VAEs) and outline our model design.
Variational Autoencoders (VAEs) are latent-variable models consisting of an encoder, a probabilistic bottleneck, and a decoder (D. P. Kingma & M. Welling, 2022b). The encoder approximates the posterior over latent variables, and the decoder reconstructs data from sampled latents. Training balances reconstruction accuracy with a regularization term enforcing a prior. To improve interpretability, $\beta$-VAE introduces a tunable weight on the regularization term for more disentangled representations (C. P. Burgess et al., 2018). $\beta$-TCVAE further decomposes the regularization into three components to encourage decorrelated probability distribution among the latents, as this has been found to promote their semantic disentanglement (R. T. Q. Chen et al., 2019).
We develop a conditional $\beta$-TCVAE model with the Convolutional Neural Network (CNN) to reconstruct the power spectra ratio $T^{2}(k,a)$, aiming to isolate the baryonic effect within the latent space, while minimizing the influence of cosmological parameters. Simultaneously, we aim for a disentangled latent space. This entails not just ensuring the latent variables are statistically independent, but also that each latent variable captures a unique feedback mechanism pattern. For this, the model is conditioned on five cosmological parameters  [^2] The total matter density $\Omega_{\mathrm{m}}$, the amplitude of matter fluctuations $\sigma_{8}$, the baryon density $\Omega_{\mathrm{b}}$, the spectral index of the primordial power spectrum $n_{\mathrm{s}}$, and the dimensionless Hubble parameter $h$. The top panel of Figure 1 shows the scheme of the model.
To obtain a redshift-independent latent, we consider three $T^{2}$ of snapshots from the same simulation ($a\approx\tfrac{1}{3},\tfrac{1}{2},1$) as one single input.
The encoder takes $T^{2}$ with the five cosmological parameters and outputs latent distribution:
$\bm{z}\sim q\bigl(\bm{z}\mid T^{2}(k,a),{\Omega_{\mathrm{m}},\Omega_{\mathrm{b}},\sigma_{8},n_{\mathrm{s}},h}\bigr),$ (2)
with scale factor $a=\tfrac{1}{3},\tfrac{1}{2},1$.
The decoder then reconstructs $T^{\prime 2}$ from the latent and cosmological parameters:
${T^{\prime}}^{2}(k,a)\sim p\bigl({T^{\prime}}^{2}(k,a)\mid\bm{z},{\Omega_{\mathrm{m}},\Omega_{\mathrm{b}},\sigma_{8},n_{\mathrm{s}},h}\bigr).$ (3)
Our evidence lower bound (ELBO) loss function comprises four terms, designed to ensure both high-quality reconstructions and informative and disentangled latent representation:
  • Reconstruction loss: Measures the quality of the model’s reconstruction of the input data.
  • Mutual information loss (MI-loss): The first KL term, corresponding to the mutual information between the latent variables and the samples.
  • Total correlation loss (TC-loss): The second KL term, promoting a disentangled latent representation. If this term vanishes, the latent distribution factorizes over its 1D marginals.
  • Dimension-wise KL loss (dw-KL-loss): Constrains each latent dimension to follow the prior distribution, typically a standard normal.
Readers can find more details in Appendix C.
II.3 Training strategy
We use the Python package Optuna to optimize the hyperparameters of the model (T. Akiba et al., 2019). Details about hyperparameter tuning and model selection are described in Appendix F.
We trained five different models. Four of them are single-dataset models, trained on IllustrisTNG SB28, Astrid SB7, Swift-EAGLE LH6 and SIMBA LH6, as mentioned in Section II.1. The fifth model is a model trained on a general dataset that includes IllustrisTNG, Astrid, and Swift-EAGLE, denoted as “TEA” in the following part of this paper. SIMBA was not included in the general dataset as its feedback model is too strong compared with the other three datasets (readers can find more about this in Appendix G). We do our test on all five different kinds of models to see if the latent space is consistent across different datasets and models. The TEA model is trained on the three datasets with the different cosmology and feedback models, so it should be able to learn a more general representation of the baryonic feedback effects.
III Results
III.1 Reconstruction
Figure 1: Top: Architecture of the conditional $\beta$-TCVAE. The encoder takes the baryonic feedback transfer function $\ln T^{2}$ and five cosmological parameters to infer the latent distribution and latent sample $\bm{z}$. Latents and cosmology are passed to the decoder to reconstruct $\ln{T^{\prime}}^{2}$. The architecture of the encoder and decoder can be found in Appendix E.
Mid left: Effect of latent dimensions on $T^{2}$ reconstruction at $z=2$, 11, and 0 in IllustrisTNG. Black points and error bars show CV-set means and variances. Solid lines show reconstruction at the latent mean, with shading for latent deviation.
Mid right: Latent PDF contours for the TEA model on IllustrisTNG, indicating disentanglement with a smooth distribution nearly factorizable along the principal axes.
Bottom: Pearson cross-correlation between latents and parameters of IllustrisTNG SB28 simulation, computed from mean latents and parameters in the validation set. Colors indicate correlation strength. Parameters are listed in Table 3.
Figure 2: Reconstruction test showing $\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}$ of the transfer functions $T^{2}$ as a function of wavenumber, averaged over redshift. For each simulation in the CV set, 100 distinct reconstructions are created. The mean square error is calculated by averaging over the whole $27\times 100=2700$ reconstructions. The redshift-averaged ratio was taken for each curve. Values near 1 indicate that reconstruction scatter is dominated by cosmic variance, with minimal additional model uncertainty. Subplots correspond to models trained on different datasets (Section II.3). Colored lines represent test sets: IllustrisTNG (blue), Astrid (orange), SIMBA (green), and EAGLE (red). Each model achieves high reconstruction accuracy on its training suite.
To verify that our model captures baryonic effects on $T^{2}$, we first assess its reconstruction performance. As we detail in Appendix C.2, the reconstructions produced by our model exhibit inherent uncertainty quantified by the latent posterior. Concurrently, our dataset also features intrinsic scatter attributed to cosmic variance. To assess if the model’s reconstruction can meet the noise levels associated with cosmic variance, we utilize the CV dataset from CAMELS introduced in Section II.1. The reconstruction of our model aligns closely with the CV within the one-sigma range as shown in the middle left panel of Figure 1 .
We further quantify this by the ratio of reconstruction MSE to cosmic variance:
$\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}=\frac{\mathrm{MSE}\left[T^{\prime 2}(k,z|T^{2}_{\textsc{cv}}),T^{2}_{\textsc{cv}}(k,z)\right]}{\mathrm{Var}\left[T^{2}_{\textsc{cv}}(k,z)\right]},$ (4)
where $T^{2}_{\textsc{cv}}(k,z)$ represents the transfer functions of CV set, and $T^{\prime 2}(k,z\mid T^{2}_{\textsc{cv}})$ are reconstructions of the model using all $T^{2}$ of the CV set as input. This metric measures reconstruction uncertainty owing to the latent scatter relative to the cosmic variance. In the ideal case that VAE learns the baryonic representation perfectly, the reconstruction is still subject to cosmic variance. Therefore, this ratio should roughly be greater than 1, and is smaller with better reconstruction.
The outcomes are shown in Figure 2 for the five models trained across various datasets. The ratio is elevated on large scales due to tiny cosmic variance down to $10^{-6}$ in that range, and progressively increases towards the small-scale end as short noise intensifies. Each model performs well on its training suite, though cross-suite performance varies. The SIMBA-trained model performs well only on SIMBA due to its strong feedback.
III.2 Latent space structure
With the reconstruction performance established in Section III.1, we now examine the latent representation. The mid-right panel of Figure 1 shows the posterior $q(\bm{z})=\sum_{\bm{x}_{n}}q(\bm{z}|\bm{x}_{n})$ for the TEA model on IllustrisTNG, revealing a smooth distribution with the two latent dimensions clearly disentangled.
This disentanglement is also evident in the reconstruction. In the mid-left panel of Figure 1, we take $\overline{T^{2}_{\textsc{cv}}}$ as input, extract the mean and standard deviation of each latent distribution, and vary one latent dimension within $\mu_{i}-0.9,\mu_{i}+0.9]$ while fixing the other at its mean. The resulting reconstructions from the decoder are shown as colored bands.
As shown in the plot, “Latent 0” enhances the spectrum suppression in the same way across all redshifts, and tends to induce feedback on large-scale. But the effect of “Latent 1” evolves with time, as a large “Latent 1” would suppress the power spectrum more strongly at $z=2$ but diminish that suppression at $z=0$. This on one hand provides additional evidence for the disentanglement of the latent space. On the other hand, the presence of two separate patterns within the latent dimensions provides hints for the physics behind them.
III.3 Correlation between latent and simulation parameters
Figure 3: Heatmaps of simulation parameter distributions projected onto the latent space. For each simulation in the IllustrisTNG suite, we generate 20 latent samples and compute the mean value of each parameter within 2D hexagonal bins. Each panel shows the average value of a specific parameter across the latent space.
From left to right, the first six panels display parameters with the strongest cross-correlations with the latent dimensions, all showing structured and consistent patterns, including detailed nonlinear features. In the right column, we present two representative parameters (a baryonic one and $\Omega_{\mathrm{m}}$) that exhibit negligible correlation with the latent space. Their heatmaps appear nearly uniform, consistent with their low cross-correlation values.
We examine how the latent parameters capture baryonic feedback by computing cross-correlations between the two latents and the 28 parameters of the IllustrisTNG SB28 simulations. For each simulation, we take the parameter values $\alpha_{i}$ and the mean of each latent’s Gaussian posterior from the encoder, yielding $28\times 2$ correlations, shown in Figure 1. Both latents show negligible correlation with cosmological parameters but distinct patterns with baryonic physics parameters, consistent with the disentanglement in Section III.2.
Of the 23 baryonic parameters, six have strong correlations (Pearson’s coefficients larger than $0.15$): (0) Wind Energy, (2) Wind Speed, (6) IMF Slope, (10) Density of Wind Recoupling, (20) BH Radiative Efficiency, and (21) Quasar Threshold. Their distributions in latent space are shown in Figure 3, with good agreement with the correlations in Figure 1. These parameters form three pairs with opposite correlation signs within each pair, suggesting complementary feedback roles:
  • (0) Wind Energy-(6) IMF slope: Both mainly correlate with “Latent 1”. High (0) Wind Energy strengthens the SN-driven winds, enhancing supernova feedback, while a steeper (6) IMF Slope has an opposite effect by increasing the metallicity gas cooling (M. E. Lee et al., 2024). This results in a strong positive correlation between SN feedback and “Latent 1”. In addition, black hole (BH) mass is also affected by (0) Wind Energy (negatively) and (6) IMF Slope (positively). Thus, “Latent 1” shows a negative correlation with BH feedback.
  • (2) Wind Speed – (10) Wind Free Travel Density Factor: Both affect wind propagation range and correlate with both latents, more strongly with “Latent 0”, influencing the scale of the feedback.
  • (20) BH Radiative Efficiency – (21) Quasar Threshold: This pair strongly correlates with “Latent 0”, and weakly with “Latent 1” in the opposite direction. Lower (20) BH Radiative Efficiency promotes BH growth, while a higher (21) Quasar Threshold favors kinetic feedback at lower accretion rates.
    As a result, larger “Latent 0” values imply stronger BH feedback, while “Latent 1” tends to suppress it.
III.4 How latents affect the spectrum
Figure 4: Upper: Heatmaps of the massive black hole fraction, $\log_{10}(f_{\mathrm{massive\ BH}})$ with $M_{\mathrm{BH}}\geq 10^{8}\,M_{\odot}$, in the latent space across $z=2$, 11, and 0 (right to left) for IllustrisTNG, similar to Figure 3. At all redshifts, “Latent 0” correlates positively with BH mass, while “Latent 1” correlates negatively.
Lower: Cross-correlation between latent variables and the massive black hole fraction across redshift for all four simulation suites. “Latent 0” shows positive correlations for IllustrisTNG and Astrid, increasing with redshift, but a negative trend for SIMBA. Negative correlations are shown with “Latent 1” for all suites.
In this section, we use the correlation patterns in Section III.3 to interpret the latent effects on $T^{2}$ reconstruction (Section III.2) and uncover their physical meaning.
Based on Section III.3, “Latent 0” is positively correlated with the BH feedback via (20) BH Radiative Efficiency – (21) Quasar Threshold, while a large “Latent 1” suppresses the BH feedback through combined effects from (0) Wind Energy, (6) IMF Slope, (20) BH Radiative Efficiency, and (21) Quasar Threshold. To explore this, we examine the fraction of massive black holes ($M_{\mathrm{BH}}\geq 10^{8}M_{\odot}$, denoted as $f_{\mathrm{massive\ BH}}$), across different redshifts in the IllustrisTNG suite, finding consistent trends between BH mass and the latents across all redshifts, as shown in the upper panel of Figure 4.
Since stronger black hole feedback suppresses the power spectrum, a large “Latent0” would enhance suppression, while a large “Latent 1” should weaken it. This is clear at $z=0$ (Figure 1, mid-left). But at $z=2$, both latents drive suppression, implying other processes beyond BH feedback are acting at early times.
SN feedback explains this discrepancy. Correlation between (0) Wind Energy–(6) IMF Slope strongly and “Latent 1” introduces SN-driven suppression, explaining the suppression at $z=2$. Meanwhile, (2) Wind Speed - (10) Wind Free Travel Density Factor correlates with “Latent 0”, yielding stronger large-scale suppression than “Latent 1”.
To test these hypotheses, we examine two special simulations: one with no BHs at all, and another with BHs but without the effective (kinetic) AGN feedback mode. In both cases, AGN feedback is effectively absent, and SN feedback dominates. We observe suppression of the power spectrum at $z=2$, supporting the idea that SN feedback, not AGN feedback, is responsible for the suppression at $z=2$. Also, similar latent values are assigned by our model: $(-1.51\pm 0.23,1.51\pm 0.63)$ for the no-BH case, and $(-1.51\pm 0.22,1.81\pm 0.63)$ for the no-kinetic-feedback case. This matches our earlier interpretation: negative Latent 0 and huge Latent 1 reflect reduced BH feedback, with big “Latent 1” coinciding with the leading SN feedback at the same time.
In summary, a large “Latent 0” corresponds to increased BH growth and stronger suppression of the power spectrum across all redshifts, especially at large scales due to farther-traveling winds. In contrast, “Latent 1” negatively correlates with BH mass and positively correlates with SN feedback, resulting in suppression at $z=2$ and diminished suppression or even enhancement at $z=0$ for large “Latent 1” values. These trends are consistent across different simulation suites, as shown by the correlation between latents and BH mass in Figure 4.
III.5 Stability of latent space dimension
Figure 5: Examples of three types of failure behavior observed in 3D latent space models.
Top left: Incomplete reconstruction from a 3D model. The red line shows the mean power spectrum ratio from the CV set, while the blue dots indicate the corresponding reconstructed spectra ratio. The reconstruction significantly deviates from the truth for $k>1\,h\,\mathrm{Mpc}^{-1}$.
Top right: An example of an overly concentrated latent distribution. The sharply striped contour indicates that at least one of the latents has collapsed into a nearly delta-function-like distribution.
Bottom: Cross-correlation between the three latent dimensions and simulation parameters similar to the bottom panel of Figure 1. For visual clarity, the sign of “Latent 1” has been flipped in this specific example to highlight its similarity to “Latent 0”.
Even though our model with a two-dimensional latent space works quite well on the general datasets, we further test whether a higher-dimensional latent space would lead to better performance. We first performed a principal component analysis (PCA) for the transfer functions. For all suites, the explained variance fractions for the first three components are $70\sim 80\%$, $10\sim 15\%$, and $5\%$, respectively, indicating that only the first two components are significant, which is consistent with early PCA results (H.-J. Huang et al., 2019). As PCA is computed directly with $T^{2}$, it contains cosmology. Thus, the number of principal components should be no less than our latent dimensionality, supporting our 2D latent space. As for the reconstruction comparison between PCA and VAE, more details can be found in Appendix I.
To confirm that two latent dimensions suffice, we further train 3D models using the same strategy as in Section II.3. However, the models on the Pareto front exhibit three types of misbehavior as shown in Figure 5: (1) incomplete reconstruction, in the top left panel, (2) an oddly behaving latent space (either overly dispersed or highly concentrated) in the top right panel, and (3) a failure to disentangle, which is shown in the bottom panel, where two latent dimensions are strongly correlated or there is one meaningless latent. The first two indicate overfitting, while the third suggests the model attempts to replicate an existing 2D latent direction rather than learning a new one. This demonstrates the model’s inability to form an independent third latent.
Combining these results with the physical interpretation of the 2D latent space in Section III.4, we conclude that two latent dimensions are sufficient to capture baryonic physics across the four CAMELS suites.
IV Discussion
In this paper, we developed a probabilistic machine learning model based on the $\beta$-TCVAE architecture to learn a general, low-dimensional latent representation of baryonic effect on matter power spectrum from multiple CAMELS simulation suites. The resulting two-dimensional latent space is both disentangled and minimally correlated with cosmological parameters, enabling simulations from different suites to be coherently mapped into a unified representation. Previous methods for modeling baryonic physics often struggle to generalize across suites and suffer degraded reconstruction accuracy from degenerate parameterizations (A. J. Zhou et al., 2025; M. Schaller & J. Schaye, 2025; M. Schaller et al., 2025; L. Kammerer et al., 2025). But our unified latent space supports seamless generalization and provides a foundation for future simulation-based inference. Furthermore, we find that the two latent dimensions modulate the transfer function $T^{2}$ primarily related to BH growth and SN feedback within the simulations we examined.
At present, our architecture shows biases when SIMBA is included in the training set, likely due to its extreme feedback model, and our physical interpretation remains largely driven by IllustrisTNG. These limitations could be alleviated with the upcoming second-generation CAMELS simulations, which feature larger volumes and more sophisticated baryonic physics. However, as shown in Figure 2, the TEA model can still reach a decent reconstruction on SIMBA within the range expected from cosmic variance. On the other hand, though we already include over $5,000$ simulations from 4 different suites, a broader and more diverse dataset would enable the model to capture more complex and extreme feedback phenomena. In parallel, more advanced architectural extensions could be explored once supported by richer training data.
Given that the two latent dimensions correlate strongly with only six baryonic parameters, it may be possible to derive an analytical mapping with symbolic regression, offering qualitative insight into feedback mechanisms. The decoder also serves as a fast emulator, rescaling gravity-only matter power spectra into their hydrodynamic counterparts. A similar approach could be extended to other summary statistics or observables, such as weak-lensing $3\times 2$pt correlation functions.
Looking ahead, the upcoming Stage-4 surveys, such as LSST, Euclid, and Roman will demand accurate modeling of baryonic effects in weak lensing analysis. With the transfer functions generated by our model, baryonic contributions can be separated from cosmology, thereby extending cosmological studies to smaller scales. By encoding spectra from real observations, our framework could directly compare feedback models against data through weak-lensing correlation functions, calibrating the latent space to real data while enabling systematic comparisons between simulations and observations.
We gratefully acknowledge the valuable comments provided by Xin Liu. Y.L. is supported by the Major Key Project of Peng Cheng Laboratory and the National Key Research and Development Program of China under grant number 2023YFA1605600. S.L. acknowledges support by Illinois Campus Research Board Award RB25035 and NSF grant AST-2308174. The Flatiron Institute is supported by the Simons Foundation. B.D. acknowledges support from the Ambrose Monell Foundation, the Corning Glass Works Foundation Fellowship Fund, and the Institute for Advanced Study. This work utilizes resources supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign (V. Kindratenko et al., 2020).
Appendix A Comparison between Methods
Table 1: Comparison across five representative methods: baryonic correction model (BCM), symbolic regression (SR), physically motivated parameterizations (PMP), principal component analysis (PCA), and our $\beta$-TCVAE method. $\checkmark$ = yes and $\times$ = no. D means data-driven, which typically uncover more parsimonious representations, and M stands for hand-crafted, that could result in large parameterizations.
Criterion BCM SR PMP PCA VAE (this work)
Driven by Data/Modeler M D M D D
Subgrid Model Independence $\times$
Cosmology Independence $\times$ $\times$
Redshift Independence $\times$
Interpretability $\times$ $\times$
Disentangled Representation $\times$ $\times$ $\times$ $\times$
As discussed in section I, numerous studies have investigated baryonic feedback effects using the transfer function. Here, we summarize the advantages and limitations of five methods: baryonic correction models (BCM), symbolic regression (SR), physically motivated parameterizations (PMP), principal component analysis (PCA), and our $\beta$ total correlation variational autoencoders ($\beta$-TCVAE) in Table 1.
Appendix B CAMELS simulations
As introduced in Section II.1, we use four suites from CAMELS: IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6. Here we summarized the basic properties of the four suites in Table 2.
Especially, the IllustrisTNG SB28 suite employs a 28-dimensional Sobol sequence to uniformly sample cosmological and astrophysical parameter space, providing broad and efficient coverage of high-dimensional variations. Each simulation utilizes a unique parameter combination and random seed, allowing for both systematic parameter studies and assessments of cosmic variance. Runs begin at $z=127$ with 2LPT initial conditions, include matched dark-matter-only counterparts, and output 91 snapshots (including $z=0,1,2$ that we use), with halos and merger trees identified using multiple standard finders. We summarize the 28 parameters in Table 3.
Table 2: Summary of the four CAMELS simulation suites used.
Suite Number Code Subgrid Model Dimension Sampling
IllustrisTNG SB28 2,048 AREPO IllustrisTNG 28D Sobol sequence
Astrid SB7 1,024 MP-Gadget ASTRID 7D Sobol sequence
SIMBA LH6 1,000 GIZMO SIMBA 6D Latin Hypercube
Swift-EAGLE LH6 1,000 Swift EAGLE 6D Latin Hypercube
Table 3: Description of parameters of IllustrisTNG SB28
Idx Parameter Idx Parameter
0 Wind Energy 1 Radio Feedback Factor
2 Wind Speed 3 Radio Feedback Reorientation
4 Max SFR Timescale 5 Factor for Softer EOS
6 IMF Slope 7 SNII Min Mass
8 Thermal Wind Fraction 9 Variable Wind Spec Momentum
10 Wind Free Travel Density Factor 11 Min Wind Speed
12 Wind Energy Reduction Factor 13 Wind Energy Reduction Metallicity
14 Wind Energy Reduction Exponent 15 Wind Dump Factor
16 Seed Black Hole Mass 17 Black Hole Accretion Factor
18 Black Hole Eddington Factor 19 Black Hole Feedback Factor
20 Black Hole Radiative Efficiency 21 Quasar Threshold
22 Quasar Threshold Power 23–27 Cosmological Parameters
Appendix C Deduction of the ELBO of VAE
C.1 The family of VAE
Variational autoencoders (VAE) are a latent variable model consisting of three parts: an “Encoder”, a “Decoder” and a “Bottleneck” in between (D. P. Kingma & M. Welling, 2022b). For a given input sample, the encoder will map the sample $\bm{x}$ to an element $\bm{z}$ in the bottleneck, called “latent” in our case, following a posterior distribution corresponds to $\bm{x}$, $q(\bm{z}|\bm{x})$. With the decoder, we could generate a new sample for a given latent, with the distribution $p(\bm{x}^{\prime}|\bm{z})$.
VAE is trained by optimizing the tractable evidence lower bound (ELBO):
$\mathscr{L}_{\mathrm{VAE}}=\frac{1}{N}\sum_{n=1}^{N}(\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]-\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))).$ (C1)
In the first term, $p(\bm{x}_{n}|\bm{z})$ stands for the likelihood of a certain sample $\bm{x}_{n}$ given the latent $\bm{z}$. If the $\bm{z}$ with a high value of $q(\bm{z}|\bm{x}_{n})$ (i.e., the latent value $\bm{z}$ preferred by the sample $\bm{x}_{n}$) also has a large $p(\bm{x}_{n}|\bm{z})$ value (i.e., likely to reconstruct the data), the first term would be large. So, the first term indicates the reconstruction power of VAE and is usually called “Reconstruction Loss”.
In the second term, $p(\bm{z})$ stands for the prior distribution of the latent, a standard Gaussian distribution is our case. When this term is small, the posterior closely matches the prior, implying that the latent representation carries little information about the input. In the extreme case where the KL divergence vanishes, the latent space has zero information capacity. Therefore, a larger KL term is often desirable to ensure informative latent encodings—this is why the ELBO is maximized despite the negative sign in front of the KL divergence term. In the extreme case where the KL divergence vanishes, the latent space has zero information capacity. Therefore, a larger KL term is often desirable to ensure an informative latent.
In practice, we would like different dimensions of the latent to be aligned with the components that have different contributions to reconstruction. So, there is betabeta-VAE, with the following “modified” ELBO:
$\mathscr{L}_{\beta}=\frac{1}{N}\sum_{n=1}^{N}\left(\mathbb{E}_{q}[\log p(\bm{x}_{n}|\bm{z})]-\beta*\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))\right).$ (C2)
With $\beta>1$, the effect of the second term would be enhanced so that a more disentangled latent may be achieved (C. P. Burgess et al., 2018).
$\beta$-TCVAE: Even though $\beta$-VAE may provide us with more disentangled latent representation, it may end up with a worse reconstruction, as much information is lost when $q(\bm{z}|\bm{x})$ approaches a standard Gaussian. To solve this problem, we may need to dig into the KL term in the ELBO. As shown in (R. T. Q. Chen et al., 2019), that term can actually be decomposed into three KL terms:
$\frac{1}{N}\sum_{n=1}^{N}(\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z})))$ (C3)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))]$ (C4)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log(q(\bm{z}|\bm{x}_{n}))-\log(p(\bm{z}))+$ (C5)
$\log(q(\bm{z}))-\log(q(\bm{z}))+\log(\prod_{j}q(\bm{z}_{j}))-\log(\prod_{j}q(\bm{z}_{j}))]]$ (C6)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|x_{n})}{q(\bm{z})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{\prod_{j}q(z_{j})}{p(\bm{z})})]$ (C7)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z},\bm{x}_{n})}{q(\bm{z})p(\bm{x}_{n})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\sum_{j}\mathbb{E}_{q(z_{j})}[\log(\frac{q(z_{j})}{p(z_{j})})]$ (C8)
$=$ $\operatorname{KL}(q(\bm{z},\bm{x}_{n})||q(\bm{z})p(\bm{x}_{n}))+\operatorname{KL}(q(\bm{z})||\prod_{j}q(z_{j}))+\sum_{j}\operatorname{KL}(q(z_{j})||p(z_{j})).$ (C9)
These are the three KL term introduced in Section II.2. As we would like to find an explainable latent representation for the baryonic feedback effects with the transfer function, we choose $\beta$-TCVAE to seek a set of disentangled latents that have desired information with feedback effects.
C.2 Reconstruction Loss
For a common Gaussian decoder (i.e., the posterior $p(\bm{x}_{n}|\bm{z})$ is a standard Gaussian), the reconstruction loss can be simplified as the Mean Squared Error (MSE) loss:
[…]
$=\frac{D}{2}\mathrm{MSE}(\bm{x}^{\prime}-\bm{x})+c,$
where $D$ is the dimensionality of the input $\bm{x}$, and $\bm{x}^{\prime}$ denotes its reconstruction. However, using a simple MSE loss can be limiting as it assumes an identity covariance matrix, which may not capture the complexity of the data distribution effectively.
To improve reconstruction fidelity, we adopt a more expressive variational decoder, where the likelihood $p(\bm{x}_{n}|\bm{z})$ is modeled as a Gaussian with a non-identity, diagonal covariance matrix:
$p(\bm{x}_{n}|\bm{z})=\mathcal{N}(\mu_{\theta}(\bm{z}),\sigma_{\theta}(\bm{z})^{2}).$ (C11)
Each dimension $i$ of the data (corresponding to different redshifts and wavevectors $k$ of the input spectrum ratio) is assigned a unique variance parameter. Following the $\sigma$-VAE design proposed in O. Rybkin et al. (2021), we set the variance based on the input and reconstruction:
$\sigma_{\theta,i}^{2}=\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i}),$ (C12)
and the reconstruction loss becomes:
$\mathbb{E}_{q(\bm{z}|\bm{x}_{n})p(\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]=-\sum_{i}\left(\log\sigma_{\theta,i}+\frac{1}{2\sigma_{\theta,i}^{2}}\,\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i})\right)$ (C13)
When training, gradient flow through $\sigma_{\theta,i}$ is stopped during backpropagation. As a result, only $\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i})$ in the second term contribute gradient, and the variance we add would only serve as a weight. This formulation enables the model to adaptively weigh reconstruction errors, allowing for a more flexible and accurate representation of the data.
Combining the results from Equation C9 and Equation C13, the ELBO for $\beta$-TCVAE can be written as:
$\mathscr{L}_{\beta-\mathrm{TC}}=$ $-\sum_{i}\Bigl[\log\sigma_{\theta,i}+\frac{1}{2\sigma_{\theta,i}^{2}}\,\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i})\Bigr]$
$-[\alpha\operatorname{KL}\bigl(q(\bm{z},\bm{x}_{n})\mathrel{\|}q(\bm{z})p(\bm{x}_{n})\bigr)$
$+\beta\operatorname{KL}\bigl(q(\bm{z})\mathrel{\|}\operatorname{{\textstyle\prod}}_{j}q(z_{j})\bigr)$
$+\gamma\sum_{j}\operatorname{KL}\bigl(q(z_{j})\mathrel{\|}p(z_{j})\bigr)].$ (C14)
Equation C14 generalizes VAE, which is a special case when $\alpha=\beta=\gamma=1$. And if $\alpha=\beta=\gamma>1$, it becomes $\beta$-VAE. Appendix D details the loss computation methods.
Appendix D Calculation of the Losses
D.1 Sampling in Minibatch
When training and tuning the model, we usually deal with a minibatch instead of the whole dataset. So, we would need a method to estimate the value of $q(\bm{z})$ with minibatch $\hat{B}_{M}$. With the “Minibatch Stratified Sampling” (MSS) method (R. T. Q. Chen et al., 2019), it has been shown that an unbiased estimator exists:
$f(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{1}{M-1}\sum_{m=1}^{M-2}q(\bm{z}|\bm{x}_{m})+\frac{N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{M-1}),$ (D1)
where $M$ is the batch size and $\bm{z}$ is originally sampled from $q(\bm{z}|\bm{x}_{n*})$. For any given $\bm{z}$ and nn*, if we sum over all the possible combination of $\bm{x}_{M-1}$ and take the average, then each $q(\bm{z}|\bm{x}_{n})$ would appear in the sum $\frac{1}{M-1}\sum_{m=1}^{M-2}q(\bm{z}|\bm{x}_{m})$ for (M2)(M-2) times and appear as $\frac{N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{i})$ once. So, for each $\bm{x}_{n}$, the total contribution would be:
$\frac{M-2}{M-1}q(\bm{z}|\bm{x}_{n})+\frac{N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{n})$ (D2)
$=\frac{N(M-2)+N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{n})$
$=\frac{N(M-1)-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{n})$
$=\frac{N-1}{N}q(\bm{z}|\bm{x}_{n}).$
Averaging over $\bm{x}_{1}$$\bm{x}_{(M-1)}$}, we would have a new estimator:
$\bar{f}(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{N-1}{N(M-1)}\sum_{m=1}^{M-1}q(\bm{z}|\bm{x}_{m}).$ (D3)
It can be easily seen that Equation D3 returns to the definition of $q(\bm{z})$ (i.e., $q(\bm{z})=\frac{1}{N}\sum_{n}q(\bm{z}|\bm{x}_{n})$) when $M=N$.
D.2 Calculation of KL terms
For each $\bm{x}_{n}$ from the minibatch $\hat{B}_{M}$, we denote the corresponding latent sample as $\bm{z}_{n}$, which follows $q(z|x)$. We label the minibatch of $\bm{z}_{n}$ as $q(\hat{B}_{M})$. $q(\hat{B}_{M})$ can also be equivalently considered as sampling by $q(\bm{z})$ from the whole parameter space. For any $\bm{z}_{n}$, we can estimate the value of $q(\bm{z}_{n})$ by the estimator from the last section.
So the value of MI-loss can be estimated by:
$\mathscr{L}_{\mathrm{MI}}$ $=\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|\bm{x})}{q(\bm{z})})]$ (D4)
$=\overline{\ln q(\bm{z}_{n}|\bm{x}_{n})-\ln q(\bm{z}_{n})}$
Here, the overline indicates averaging over the whole $q(\hat{B}_{M})$. As $q(\hat{B}_{M})$ is constructed by sampling with $q(\bm{z}|\bm{x})$, averaging over it can be considered an estimator of the expectation.
Similarly, we can estimate the TC-loss in the same way:
$\mathscr{L}_{\mathrm{TC}}$ $=\int q(\bm{z})\ln\left(\frac{q(\bm{z})}{\prod_{j}q(z_{j})}\right)\mathrm{d}z$ (D5)
$=\overline{\ln q(\bm{z}_{n})-\ln(\prod_{j}q(z_{n,j}))}.$
In practice, as we would like a set of disentangled parameters, $\mathscr{L}_{\mathrm{TC}}$ would be close to 0. In this case, as $q(\bm{z}_{n})$ and $\prod_{j}q(z_{n,j})$ are not perfectly normalized due to sampling, the difference in their normalization may result in a negative $\mathscr{L}_{\mathrm{TC}}$. However, adding the normalization factor into the estimator would overestimate the TC loss. As a result, we decided to keep the unbiased form of Equation D5 while keeping in mind that a small negative $\mathscr{L}_{\mathrm{TC}}$ might occur.
D.3 Cyclical Annealing
A problem with both $\beta$-VAE and $\beta$-TCVAE is that, at the beginning of training process, the latent $\bm{z}$ can hardly represent the dataset due to the random initialization. However, all those KL terms in the ELBO would push the posterior $q(\bm{z}|\bm{x})$ to the uninformative prior $p(\bm{z})$. As a result, the decoder would tend to do reconstruction while ignoring the latents, which is usually called “Latent Space Collapse”.
To deal with this problem, we choose to do cyclical annealing during training.H. Fu et al. (2019) To be specific, we add a new parameter $\lambda$ to control all the KL terms in the ELBO and periodically varying it in the whole training process:
$\mathscr{L}^{\prime}_{\beta-\mathrm{TC}}$ (D6)
$=$ $\mathbb{E}_{q(\bm{z}|\bm{x}_{n})p(\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]$
$-\lambda[\alpha\operatorname{KL}(q(\bm{z},\bm{x}_{n})||q(\bm{z})p(\bm{x}_{n}))+\beta\operatorname{KL}(q(\bm{z})||\prod_{j}q(z_{j}))+\gamma\sum_{j}\operatorname{KL}(q(z_{j})||p(z_{j}))].$
When $\lambda$ is small, the ELBO would be dominated by the reconstruction loss, so the model will focus on learning to accurately reconstruct the data, ensuring that the latent variables capture significant information about the inputs. On the other hand, during the high $\lambda$ phase, KL terms becomes important, which helps in learning a disentangled and structured latent space. By alternating between these phases, cyclical annealing allows the model to find a better balance, improving both reconstruction quality and latent space disentanglement without falling into the extremes of latent space collapse or under-regularization.
Appendix E Model Architecture
We utilize a combination of multilayer perceptron (MLP) architecture and pointwise convolution layers to construct our Variational Autoencoder (VAE) model, ensuring a symmetric and efficient structure for encoding and decoding.
The encoder is designed to process $T^{2}(k,z)$ from three snapshots along with the five cosmological parameters. For the first step, a pointwise convolution is applied to each k-bin independently. This operation increases the feature dimensionality at each bin without changing the number of $k$-bins, allowing the model to extract richer local representations. The resulting intermediate features are then concatenated with the five cosmological parameters and passed into a multi-layer perceptron (MLP). The MLP then output the mean and variance of the latent variables, forming the posterior distribution.
The decoder takes the sampled latent variables together with the same five cosmological parameters as input. These inputs are first processed by another MLP, which produces an intermediate representation. This intermediate output is reshaped to match the number of $k$-bins in the power spectrum and then passed through a pointwise convolution layer to reconstruct $T^{2}$. To incorporate the variational decoder as mentioned in Appendix C.2, we also output the variance of the reconstructed $T^{2}$, $\sigma_{\theta,j}=\mathrm{MSE}(\bm{x}^{\prime}_{j},\bm{x}_{j})$. This approach allows the model to learn a more flexible and accurate representation of the data by considering the variance in the reconstruction process (O. Rybkin et al., 2021).
Table 4: All Hyperparameters Tuned by Optuna. The range for parameters is just for reference. For a specific training run, the range would be changed, and some parameters might be fixed for a model with the desired behavior. For ord, “B”, “A”, and “D” stand for “Batchnorm”, “Activation function”, and “Dropout”. lgta is the leaky slope transformed with the Sigmoid function. For mu and nu, the number of epochs in each annealing cycle is $1/(\texttt{mu}\cdot\texttt{nu})$ and the duration of annealing is $1/\texttt{mu}$.
Category Name Type Range Description
Model Dc int [1, 10] Depth of convolutional layers
2Wc int [2, 10] Convolution layer width is $2^{\texttt{2Wc}}$
Dm int [1, 10] Depth of MLP
2Wm int [4, 10] MLP layer width is $2^{\texttt{2Wm}}$
lgta float [-5, 5] Logit of leaky ReLU slope
ord categorical {BA, A, AB, Module order in architecture
BAD, AD, ABD}
Loss lambda float (log) $10^{-4},10^{-1}]$ Overall KL scaling
mu float (log) $2/\texttt{num\_epoch},1]$ Controls annealing duration
nu float (log) $[0.1,1]$ Controls annealing frequency
alpha float (log) $[0.1,10]$ Weight for MI loss
beta float (log) $[0.1,1]$ Weight for TC loss
gamma float (log) $10^{-3},10]$ Weight for dw-KL loss
Training lr float (log) $10^{-6},10^{-1}]$ Learning rate
1-b1 float (log) $10^{-3},1]$ $1-\beta_{1}$ for AdamW
1-b2 float (log) $10^{-6},1]$ $1-\beta_{2}$ for AdamW
wd float (log) $10^{-6},10^{-1}]$ Weight decay
inid categorical {uniform, normal} Initialization distribution
inim categorical {fan_in, fan_out} Initialization mode
bnmf float (log) $10^{-2},10^{2}]$ BN momentum
dopc float (log) $[0.01,0.5]$ Dropout rate in CNN
dopm float (log) $[0.01,0.5]$ Dropout rate in MLP
Figure 6: An example of the training and validation loss curves. We use a learning rate scheduler with a patience of 1000 epochs and a 1% threshold, along with early stopping if no improvement is observed within 2000 epochs. Due to the cyclic annealing strategy described in Appendix D.3, the KL and reconstruction losses may exhibit oscillations during training, but they generally converge to stable values by the end.
Figure 7: Pareto front of TEA models selected during hyperparameter optimization. Each point represents a trial in the OPTUNA search space, plotted by its reconstruction loss (horizontal axis) and total KL loss (vertical axis), with color indicating the total correlation (TC) loss. Gray circles denote all trials, colored squares highlight the Pareto-optimal solutions, and the blue star marks the final model selected for all the plots shown in the paper.
Appendix F Model Tuning
As noted in Section II.3, we optimize 21 hyperparameters related to the model architecture, training procedure, and loss formulation using the Python package OPTUNA. A complete list of these hyperparameters is provided in Table 4. In addition to the optimized settings, we fix the batch size to match the size of the training (or validation) set to avoid bias in loss estimation that can arise from small batches.
OPTUNA searches for optimal hyperparameter configurations by simultaneously minimizing the reconstruction loss, KL divergence loss, and total correlation (TC) loss. To avoid selecting models with degenerate latent posteriors, we impose the following constraint:
$1\leq A=\frac{\mathrm{Var}(\mu(\bm{x}))}{\overline{\sigma(\bm{x})^{2}}}\leq 10,$ (F1)
where $\mu(\bm{x})$ and $\sigma(\bm{x})$ denote the mean and standard deviation of the latent posterior for a given input $\bm{x}$, and the variance is computed across the entire dataset. This criterion ensures that the latent representation is neither too concentrated (i.e., collapsing to a nearly constant value) nor overly dispersed (i.e., resembling delta functions as in traditional autoencoders). All models that fail to satisfy this condition are pruned, and only the remaining models are considered in the later analysis.
Due to the inherent trade-off between reconstruction loss and regularization losses (KL and TC), the hyperparameter optimization yields a Pareto front rather than a single global minimum. As illustrated in Figure 7, most Pareto-optimal models lie along a frontier governed primarily by reconstruction and KL loss, with a few outliers exhibiting exceptionally low TC loss (depicted as blue squares). With the Pareto-optimized parameters, we retrain the models using those hyperparameters with a longer training process, and then select the best models among the retrained ones. An example of the retraining loss curve is shown in Figure 6. The final model is selected by balancing reconstruction accuracy with latent disentanglement, and is highlighted by a star in the plot.
Appendix G Results with SIMBA
Figure 8: Same as Figure 2, but for the EAST and “TEA tuned” models. While the EAST model performs well on the SIMBA suite, it exhibits significantly higher variance on the other suites, particularly at low-$k$. In contrast, the “TEA tuned” model achieves consistently better reconstruction across all suites.
Figure 9: Same as the middle-left panel of Figure 1, but showing results from the EAST and “TEA tuned” models on the Astrid suite.
Figure 10: Range of $T^{2}$ for different suites. The blue region stands for the range between upper and lower limit in each $k$ bin. The red region represents the $16-84\%$ range in each $k$ bin. Comparing to the other suites, SIMBA both significantly deviate from 1 and have much larger variance on the largest scales.
As discussed in Section II.3, our general training dataset for the TEA model does not include SIMBA LH6 due to its strong baryonic feedback. In this section, we present results involving the SIMBA suite to further justify the construction of the “TEA” model.
We compare two models. The first, denoted as EAST, is trained on all four simulation suites (IllustrisTNG, Astrid, SIMBA, and EAGLE) using the same hyperparameter setup described in Appendix F. The second, called the “TEA tuned” model, starts from the EAST model and is further fine-tuned on the TEA dataset, which excludes SIMBA.
Figure 8 shows reconstruction performance on the cross-validation set. The EAST model achieves excellent reconstruction on SIMBA but performs poorly on the other three suites, especially at large scales (low-$k$), where the variance is much larger. In contrast, the “TEA tuned” model exhibits significantly improved performance on the other suites while maintaining acceptable behavior on SIMBA.
This issue is further illustrated in Figure 9, which mirrors the latent variation analysis shown in the middle-left panel of Figure 1. The EAST model shows strong reconstruction bias at low $k$ in the Astrid suite, and the latent dimensions exert a large influence in that regime. After fine-tuning on the TEA dataset, the “TEA tuned” model significantly reduces this bias, particularly in the large-scale regime. A reason for this is that, SIMBA not only have a stronger feedback than the other suite on the largest scales, but also have larger variance due to the wide range of CAMELS, as shown in Figure 10. Based on these results, we eventually decide to use TEA model for the main analysis shown in this paper.
Appendix H Reproducibility Test
Figure 11: Cross-correlation matrix between latent space from 10 independently trained TEA models, where “M0” is the one we choose for the main result of the paper. Each model is trained with identical hyperparameters on combined data from four simulation suites (IllustrisTNG, Astrid, SIMBA, and EAGLE). To account for the inherent symmetry in the latent space, we align each model pair by evaluating all possible axis permutations and sign flips, reporting the maximum average absolute Pearson correlation across the two latent dimensions. All model pairs exhibit strong correlations ($\geq 0.6$), with a mean of 0.84, indicating robust reproducibility of the learned latent space despite stochastic variations in training.
Due to the stochastic nature of the VAE training process, we assess reproducibility by training 10 models with identical hyperparameters and analyzing the cross-correlation between their latent representations. In Figure 11, we present the average absolute cross-correlation across all 10 models, evaluated on the combined data from the four simulation suites. The “M0” model is the “TEA” model mentioned in the main text. To account for the inherent symmetry in latent space, we consider all possible axis permutations and sign flips, and report the maximum correlation for each model pair. All model pairs exhibit cross-correlations above 0.6, with a mean value of 0.84, demonstrating a high degree of consistency across retrainings and the robustness of our latent space.
Appendix I PCA Reconstruction Test
Figure 12: Reconstruction RMSE of matter power spectrum ratios on the test set for our VAE models, compared with PCA using varying numbers of principal components (PCs). Each panel corresponds to one simulation suite: IllustrisTNG, Astrid, SIMBA, EAGLE, and a combined “TEA” training case. Here PCA is trained on the same training data as the corresponding VAE. In the “TEA” panel, PCA is trained on the TEA training set for comparison to the reconstruction of the TEA model, and we show reconstruction performance on the joint test set across the three suites of TEA (IllustrisTNG, EAGLE, and Astrid).
Solid curves show PCA reconstruction errors for $n_{\rm pc}=1\sim 5$, with darker colors indicating fewer components, while the blue dashed curve shows the VAE reconstruction. The horizontal dashed line marks the 1% RMSE threshold.
Across all scales, our 2-dimensional VAE achieves better reconstruction than PCA with 2 PCs and performs comparably to PCA with 3 PCs. Notably, the VAE outperforms PCA with 5 PCs on large scales for all suites. Moreover, around $k=1\,h/\mathrm{Mpc}$, PCA reconstruction rarely reaches the 1% RMSE level even with 5 PCs, while the VAE meets or approaches this accuracy for all cases except for SIMBA.
To evaluate the capability of our VAE models versus the conventional linear principal component analysis (PCA) approach (T. Eifler et al., 2015; H.-J. Huang et al., 2019), we compare their reconstruction root-mean-square errors (RMSE) across simulation suites. The RMSE is defined as: $\mathrm{RMSE}(T^{2})=\sqrt{\overline{(T^{\prime 2}-T^{2})^{2}}}$.
For a fair comparison, PCA is trained using the same training data as the corresponding VAE model for each suite (IllustrisTNG, Astrid, EAGLE, SIMBA, and TEA). The principal components are obtained by diagonalizing the covariance matrix of $T^{2}$. The resulting reconstruction errors are shown in Figure 12.
As illustrated in the figure, our 2-dimensional VAE models consistently outperform PCA with two PCs and achieve comparable performance to PCA with three PCs. On large scales, the VAE even surpasses PCA with five PCs across all suites. In particular, at $k\approx 1\,h/\mathrm{Mpc}$, the VAE reconstructs the transfer function with $\sim$1% RMSE, whereas PCA fails to reach this accuracy in all suites.
This performance gap is expected because PCA captures only the leading linear modes of variation in the transfer function, while the VAE learns a nonlinear latent representation. As a result, PCA requires multiple PCs to reach the same accuracy at the same $k$ and has to sacrifice low $k$’s and likely some parameter choices, whereas the VAE can efficiently encode and reconstruct these effects using only two latent variables.